OpenStudy (anonymous):

if h varies directly as d and inversely as f, and h=24 when d=12 and f=5, find h when d=7 and f=9

OpenStudy (dumbcow):

h =kd/f 24 = 12k/5 120 = 12k 10=k -> h = 10d/f h = (10*7)/9 h = 70/9